# Are foundation models for computer vision good conformal predictors?

### v2

@@ Abstract
Recent advances in self-supervision and contrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has been barely explored. In this work, we delve into the behaviour of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. We also show that calibrating the confidence predictions of these models, a popular strategy to improve their uncertainty quantification, actually leads to efficiency degradation of the conformal set on adaptive CP methods. Furthermore, few-shot adaptation of Vision-Language Models (VLMs) to downstream tasks, whose popularity is surging, enhances conformal scores compared to zero-shot predictions. Last, our empirical study exposes APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage guarantees across multiple challenging, yet realistic scenarios.
@@ 1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 [40], as well as those integrating text, such as CLIP [46], are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out to the existence of bias and factual errors in these models [58], which transcend the field of computer vision [52, 4]. For example, the original CLIP paper [46] demonstrated gender and race biases in certain zero-shot tasks, whereas [58] identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet models. Furthermore, [36] recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare or security applications.
@@
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) [16], a simple variant of Platts Scaling [44], provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other lines of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space [45, 43], logits [32, 31], or modifying the ground truth labels [34, 35].
@@
Conformal Prediction (CP) [60], an alternative strategy to quantify the uncertainty [49, 48, 1], is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions, contrary to calibration approaches.
@@
Due to these properties, CP is gaining attention to conformalize the predictions of deep models [2, 1, 11, 50]. Nevertheless, albeit the efforts to study CP in large language foundation models [15], its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on eleven popular vision foundation models across multiple vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
@@
    Vision, and vision-language foundation models seem to better conformalize compared to their more traditionally (fully-supervised) trained counterparts.
@@
    Across different foundation models, those including visual transformers, such as DINO and CLIP, lead to better conformal results compared to models integrating convolutional neural networks.
@@
    Under distributional shifts, APS exhibits the highest robustness among CP methods in terms of coverage guarantes, albeit decreasing its set efficiency.
@@
    Confidence calibration decreases the efficiency of conformal sets, but typically improves coverage gap.
@@
    Few-shot adaptation of vision-language models (VLMs) yields better conformalization than zero-shot predictions in ID data, with marginal gains on OOD.
@@ 2 Related Work
Foundation models for computer vision.  The landscape of foundation models has rapidly evolved in the last years. Traditionally, pre-trained convolutional networks based on ResNet architectures [18] were the main models used by the community. However, driven by the unprecedented advances in language models, e.g., GPT [7] or LLaMA [57], as well as the vast availability of image data online, there are groundbreaking advances in unimodal [8, 40, 26] and multimodal [46] foundation models for vision tasks, commonly based on vision transformers. These large pre-trained models aim to generalize across a broad span of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models such as DINO [8, 40] rely on self-supervised learning strategies on large datasets, leading to excellent semantic understanding of visual content. On the other hand, CLIP [46] bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
@@
Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches [16, 12, 24], which map the logits or softmax predictions to smoother distributions, and explicit learning objectives [31, 32, 35, 34, 19, 37], which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples [61, 51]. Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different” a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a user-specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class conditional coverage. For this purpose, several non-conformity scores have been presented [56, 48, 1, 49, 13, 11, 55], with [48, 1, 49] being popular methods widely studied. A straightforward solution directly uses the raw class softmax predictions to generate the prediction sets [49]. Adaptive Prediction Sets (APS) [48] provides an adaptive version, computing non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficiency, RAPS [1] introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
@@
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on data collection that falls within the calibration and test data points distribution. Despite this transfer learning framework not necessarily affecting the marginal guarantees provided by CP, how it affects its efficiency and conditional coverage remains to be explored. Thus, quantifying the uncertainty of their predictions is paramount given the rising popularity of foundation models in strategic domains. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied [36, 66, 59], its exploration under CP is, to our knowledge, overlooked.
@@ 3.1 Conformal Prediction Framework
Coverage Guarantees.  A key property of conformal prediction is its finite-sample coverage guarantee. This property ensures that the prediction sets achieve the desired coverage probability marginally over $\mathcal{X}$ and $\mathcal{Y}$, irrespective of the underlying data distribution, as long as the calibration and test data are exchangeable [61]. Formally, for any $1-\alpha$, conformal predictors satisfy:
@@ 3.2 Non-conformity scores
Least Ambiguous Classifier (LAC) [49]  aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
$\mathcal{S}_{\text{\tiny{LAC}}}(\mathbf{x},y)=1-x_{k=y}.$ (4)
LAC also provides notable efficiency in scenarios using an imperfect classifier. However, it lacks adaptability, e.g., in under-represented categories or uncertain predictions.
Adaptive Prediction Sets.  APS [48] provides a non-conformity score that leverages the accumulated confidence in the ordered probability predictions. Thus, APS is known to be an adaptive score, whose main objective is enhancing the coverage of uncertain predictions by sacrificing efficiency. Formally, APS is expressed as:
$\mathcal{S}_{\text{\tiny{APS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u,$ (5)
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to archieve exact marginal coverage.
Regularized Adaptive Prediction Sets.  RAPS [1] builds upon APS by adding a regularization term to enforce smaller predicted sets. Thus, APS score is modified to penalize the confidence of introducing additional, unlikely categories, after a certain set size is met:
$\mathcal{S}_{\text{\tiny{RAPS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u+\lambda(o(\mathbf{x},y)-k_{reg})^{+}$ (6)
@@
We refer the reader to the different works [1, 49, 48] for the respective conformal calibration coverage guarantees.
@@ 4.1 Experimental Setup
Models: We employ a total of 11 foundation models: two DINO [8] (DINO-S and DINO-B), four DINOv2 [40] (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg [3] (with ResNet-50, ResNet-50x2, and ResNet-200x2), and two VLMs (CLIP [46] and MetaCLIP [65]) based on ViT-B. Our main analysis is conducted on three popular vision datasets: CIFAR-10 [28], CIFAR-100 [28], and Imagenet [10], including its versions integrating domain shifts [21, 22, 62, 47]. Each dataset is split into two sets: one for training, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation. For few-shot, we adhere to the emerging CLIP few-shot literature [68, 53, 69], and evaluate models on 10 additional fine-grained and general concepts classification benchmarks: SUN397 [64], FGVCAircraft [33], EuroSAT [20], StanfordCars [27], Food101 [6], OxfordPets [41], Flowers102 [39], Caltech101 [14], DTD [9], and UCF101 [54].
Metrics: We resort to conformal set size (a.k.a. efficiency), minimum class-conditional coverage (MCCC), coverage gap (CovGap) and marginal coverage to benchmark the different methods. More details about metrics and implementation details can be found in Appendix A and Appendix B.
Adaptation to target tasks.  The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do not accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen, and a linear probing (LP) head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss (more details in Appendix B).

### v3

@@ Abstract
Recent advances in self-supervision and contrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has received little attention. In this work, we delve into the behaviour of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. We also show that calibrating the confidence predictions of these models, a popular strategy to improve their uncertainty quantification, actually leads to efficiency degradation of the conformal set on adaptive CP methods. Furthermore, few-shot adaptation of Vision-Language Models (VLMs) to downstream tasks, whose popularity is surging, enhances conformal scores compared to zero-shot predictions. Last, our empirical study exposes APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage guarantees across multiple challenging, yet realistic scenarios.
@@ 1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 (Oquab et al., 2024), as well as those integrating text, such as CLIP (Radford et al., 2021), are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out the existence of bias and factual errors in these models (Tu et al., 2023), which transcend the field of computer vision (Shuster et al., 2022; Barocas et al., 2023). For example, the original CLIP paper (Radford et al., 2021) demonstrated gender and race biases in certain zero-shot tasks, whereas Tu et al. (2023) identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet-trained models. Furthermore, Murugesan et al. (2024a) recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare or security applications.
@@
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) (Guo et al., 2017), a simple variant of Platt Scaling (Platt et al., 1999), provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other lines of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space (Popordanoska et al., 2022; Pereyra et al., 2017), logits (Liu et al., 2023a; 2022), or modifying the ground truth labels (Mukhoti et al., 2020; Müller et al., 2019).
@@
Conformal Prediction (CP) (Vovk et al., 1999), an alternative strategy to quantify the uncertainty (Sadinle et al., 2018; Romano et al., 2020; Angelopoulos et al., 2020), is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions, contrary to calibration approaches.
@@
Due to these properties, CP is gaining attention to conformalize the predictions of deep models (Barber et al., 2023; Angelopoulos et al., 2020; Ding et al., 2024; Sesia & Romano, 2021). Nevertheless, albeit the efforts to study CP in large language foundation models (Gui et al., 2024), its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on 17 popular vision foundation models across multiple vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
@@
    Vision, and vision-language foundation models seem to lead to lower set sizes and higher class-conditional coverage compared to their more traditionally (fully-supervised) trained counterparts.
@@
    Under distributional shifts, APS exhibits the highest robustness among CP methods in terms of coverage guarantees, albeit decreasing its set efficiency.
@@
    Confidence calibration decreases the efficiency of conformal sets, but typically improves coverage gap
@@
    Few-shot adaptation of vision-language models (VLMs) yields lower set sizes and coverage gaps than zero-shot predictions in ID data, with marginal gains on OOD.
@@
    Under domain shift, across different foundation models, those including visual transformers, such as DINO and CLIP, lead to smaller decreases in the conformal metrics compared to models integrating convolutional neural networks.
@@ 2 Related Work
Foundation models for computer vision. The landscape of foundation models has rapidly evolved in recent years. Traditionally, pre-trained convolutional networks based on ResNet architectures (He et al., 2016) were the main models used by the community. However, driven by the unprecedented advances in language models, e.g., GPT (Brown et al., 2020) or LLaMA (Touvron et al., 2023), as well as the vast availability of image data online, there are groundbreaking advances in unimodal (Caron et al., 2021; Oquab et al., 2024; Kirillov et al., 2023) and multimodal (Radford et al., 2021) foundation models for vision tasks, commonly based on vision transformers. These large pre-trained models aim to generalize across a broad span of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models such as DINO (Caron et al., 2021; Oquab et al., 2024) rely on self-supervised learning strategies on large datasets, leading to excellent semantic understanding of visual content. On the other hand, CLIP (Radford et al., 2021) bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
@@
Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches (Guo et al., 2017; Ding et al., 2021; Joy et al., 2023), which map the logits or softmax predictions to smoother distributions, and explicit learning objectives (Liu et al., 2022; 2023a; Müller et al., 2019; Mukhoti et al., 2020; Hebbalaguppe et al., 2022; Murugesan et al., 2024b), which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples (Vovk et al., 2005; Shafer & Vovk, 2008). Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different" a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a user-specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class-conditional coverage. For this purpose, several non-conformity scores have been presented (Stutz et al., 2022; Romano et al., 2020; Angelopoulos et al., 2020; Sadinle et al., 2018; Einbinder et al., 2022; Ding et al., 2024; Straitouri et al., 2023), with Romano et al. (2020); Angelopoulos et al. (2020); Sadinle et al. (2018) being popular methods widely studied. A straightforward solution directly uses the raw class softmax predictions to generate the prediction sets (Sadinle et al., 2018). Adaptive Prediction Sets (APS) (Romano et al., 2020) provides an adaptive version, computing non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficiency, RAPS (Angelopoulos et al., 2020) introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
@@
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on data collection that falls within the calibration and test data points distribution. Despite this transfer learning framework not necessarily affecting the marginal guarantees provided by CP, how it affects its efficiency and conditional coverage remains to be explored. Thus, quantifying the uncertainty of their predictions is paramount given the rising popularity of foundation models in strategic domains. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied (Murugesan et al., 2024a; Yoon et al., 2024; Tu et al., 2024), its exploration under CP is, to our knowledge, overlooked.
@@ 3.1 Conformal Prediction Framework
Coverage Guarantees. A key property of conformal prediction is its finite-sample coverage guarantee. This property ensures that the prediction sets achieve the desired coverage probability marginally over $\mathcal{X}$ and $\mathcal{Y}$, irrespective of the underlying data distribution, as long as the calibration and test data are exchangeable (Vovk et al., 2005). Formally, for any $1-\alpha$, conformal predictors satisfy:
@@ 3.2 Non-conformity scores
Least Ambiguous Classifier (LAC) (Sadinle et al., 2018) aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
$\mathcal{S}_{\text{\tiny{LAC}}}(\mathbf{x},y)=1-\pi_{x}(y).$ (4)
Where $\pi_{x}(y)=P(Y=y|X=x)$ is the softmax of the true class. LAC also provides notable efficiency in scenarios using an imperfect classifier. However, it lacks adaptability, e.g., in under-represented categories or uncertain predictions.
Adaptive Prediction Sets (APS) (Romano et al., 2020) provides a non-conformity score that leverages the accumulated confidence in the ordered probability predictions. Thus, APS is known to be an adaptive score, whose main objective is enhancing the coverage of uncertain predictions by sacrificing efficiency. Formally, APS is expressed as:
$\mathcal{S}_{\text{\tiny{APS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+\pi_{x}(y)\cdot u,$ (5)
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to achieve exact marginal coverage.
Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2020) builds upon APS by adding a regularization term to enforce smaller predicted sets. Thus, APS score is modified to penalize the confidence of introducing additional, unlikely categories, after a certain set size is met:
$\mathcal{S}_{\text{\tiny{RAPS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+\pi_{x}(y)\cdot u+\lambda(o(\mathbf{x},y)-k_{reg})^{+}$ (6)
@@
We refer the reader to the different works (Angelopoulos et al., 2020; Sadinle et al., 2018; Romano et al., 2020) for the respective conformal calibration coverage guarantees.
@@ 4.1 Experimental Setup
Models: We employ a total of 17 foundation models: two DINO (Caron et al., 2021) (DINO-S and DINO-B), four DINOv2 (Oquab et al., 2024) (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg (Bardes et al., 2022) (with ResNet-50, ResNet-50x2, and ResNet-200x2), and eight VLMs (five CLIP (Radford et al., 2021) models, MetaCLIP (Xu et al., 2024), LLaVa (Liu et al., 2023b), and Phi (Abdin et al., 2024)). Our main analysis is conducted on three popular vision datasets: CIFAR-10 (Krizhevsky & Hinton, 2009), CIFAR-100 (Krizhevsky & Hinton, 2009), and Imagenet (Deng et al., 2009), including its versions integrating domain shifts (Hendrycks et al., 2021a; b; Wang et al., 2019; Recht et al., 2019). Each dataset is split into two sets: one for training, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation. For few-shot, we adhere to the emerging CLIP few-shot literature (Zhang et al., 2022; Silva-Rodríguez et al., 2024; Zhou et al., 2022a), and evaluate models on 10 additional fine-grained and general concepts classification benchmarks: SUN397 (Xiao et al., 2010), FGVCAircraft (Maji et al., 2013), EuroSAT (Helber et al., 2018), StanfordCars (Krause et al., 2012), Food101 (Bossard et al., 2014), OxfordPets (Parkhi et al., 2012), Flowers102 (Nilsback & Zisserman, 2008), Caltech101 (Fei-Fei et al., 2004), DTD (Cimpoi et al., 2014), and UCF101 (Soomro et al., 2012).
[…]
$\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}|\hat{C}(\mathbf{x})|$ (7)
Where Equation 7 is the average predicted set size across the test set (sometimes referred to as efficiency), Equation 8 is the empirical marginal coverage, Equation 9 represents the average gap between the empirical coverage and the class-conditioned coverage, and Equation 10 is the min class-conditioned coverage (MCCC).
Adaptation to target tasks. The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do not accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen, and a linear probing (LP) head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss (more details in Appendix Section A).